UT 20150529 LUCI2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, APramskiy, DThompson
Remote from Tucson for training: MEdwards, OKuhn, BRothberg, RMWagner
Observer AO: N/A
Telescope Operator: SAllenson
LBTO Support: TSargent
Watchdog: RGredel (LUCI-team)

Plan for the night

- Some of the night will be used for training LBTO SA's (connected remotely from Tucson)

Summary

- Successfull training of SA's

- Stability-tests LUCI2 cont'd

- Fire (humidifier) alarm

Details

All times are UT
Files are luci2.20150529.NNNN.fits

01:40 m31rv_dark_10Nx6Dx10s.xml
0007 0016 10 x [6 x 10.0s] blind mask and filters...MER mode!!
0017 0026 10 x [6 x 10.0s] LIR/INT

02:25 Sunset, mostly clear (just a few Cirrus around), T = +9deg, Hum = 30%, low wind

HIP53735 for G210 grating throughput calibrations, also folks in Tucson practicing acquisitions

Error on the readout panel on the first image we tried taking tonight. This could have
been user error. We transferred the GEIRS gui to the Tucson computer (cleanup and
start_luci2_new) while an integration was going. Alexander had to correct something in the
SQL database (like it had forgotten what the last and next files were) and it worked after that.

02:50 Steve checks pointing and collimates

### SDSSJ1126+1913 11 26 19.42 +19 13 29.3 # PA=320 (E is left) 1.5" slit
### HIP_56736 11 37 57.81 +15 46 36.8 # 8.80 A0V
I did not note the filenumbers. Merger/AGN target for Barry

>>>Grating status panel labeling:
Currently
210 zJHK in large print and 2.20um [#2] Stable

G210 @ 2.20um all large text
Also, create a timeout watcher for grating stabilization. We don't want
a single "not stable" measurement to trigger red text on the GUI.

>>>Mask/MOS unit status panel labeling
Currently
"mask in FPU" with small type LS2.0_1200mue[4] and 4
Should be: LS2.0 in FPU or LS2.0 in Turnout
When in turnout, background yellow and text black
We will need to visit the mask naming to keep them short, or truncate

>>>Calibration unit status panel

Should say: "CU in" or "CU out"
Consider changing background of status to yellow (or blue) when CU in

>>>We will need an easy way to open the telescope service GUI, or a more user-oriented

subset of the functionality that they need in

05:29 SDSSJ134420 Ks imaging 20x5.0s LBTB_Busch
We did not run all three filters because it would have taken too long.
about 0084 - 0110 Ks images
  1. 0115 H band images (forgot to insert the pause)

>>>RTD several times slowed down while calculating a difference image, or did not

display a valid difference image. Closing and restarting the RTD did not clear
this problem. Several times I also had trouble on the LiveView page where I requested
a difference image and the result displayed was definitely not a difference of two
LIR mode images.

07:02 NGC6205_ima.xml
The script auto-started! Seems the pause I inserted in the previous script was
ignored and I loaded a new script while the detector was integrating. When the queue
went to the next item to execute, it was the first item in the new script.

Jochen making a corrected version of this script. I'll start the old version to get the
telescope collimated, and the initial data will be usable.

> Steve corrected pointing

07:05 NGC6205_ima.xml
Queue subwindow went blank! Preset had completed and the telescope service was waiting for
collimation. Guide star was likely incorrect, Steve cancelled the preset and corrected the
pointing. It was around then that the queue went blank. Alexander restarted the scheduler.

07:22 NGC6205_mod.xml

0117 0121 The first set of 1x4.0s Ks images...poorly collimated.
Did wait for collimation fail or is it something else? GCS looks OK, but images on LUCI
are really out of focus. See 0122 (taken manually after pausing script.

IQ on LUCI2 is really poor! Guide star looks OK @ EL 86 0123 before initializing detector stage
0124 after detector init...still bad
0125 FC on...still bad
Camera undefined on instrument manager gui...init camera
0126 after camera init...still bad
0127 N1.8 camera...look a little triangular 4.2pix FWHM
0128 N3.75 camera...still bad
Move in OpticSieve mask
0129 Optic Sieve mask dots look fine (<2pix to corners of detector)...its external
0130 Clear active optics...LUCI and guider images look the same (aberrated, mostly coma)
0131 after recollimating...still bad
0132 15x4.0s LIR/INT still elongated. 6x5 pixels (0.75") while guider sees 0.67"
0133 15x4.0s Slewing to a lower EL field ~64 deg (BS9158)

Go back to NGC6205_mod.xml
Images not great, but WAY better than before. If this persists, there will have to be an
Elevation limit for LUCI2.

>>>RTD again is showing the latest image as "luci2.20150529" with no sequence #

08:17 commit. Now @ 80.3d IQ still not great.
by end of second group (10th image in script, #0144) EL~79.3, images rounder
the subsequent 1x4.0s images were a little less good
  1. 1x20.0s H band images are rounder but have haloes (spherical aberration)

08:33 Steve will try clearing active optics, homing M2 hexapod, initializing mirrors,
restarted GCSr, home AGw, rest and flatten shell

08:40 Re-submit NGC6205_mod.xml
pause after first image (0154) EL~75.75
skip to longer K exposure (0155) round, but not a nice PSF
  1. 10 x 2.75s stars look better. I think the brighter stars on 13 were just saturating.
    Definitely had bad IQ at higher EL though.
    Skipping remainder of this target

08:59 PN_Vy2-2_lowres.xml
  1. 0158 0159 acq 0160 confirm
    Lots of lines!

>>>FW2 error on switching between zJspec and HKspec

FW2 init did a long series of +10 steps, ended at Free11
Manually moved to HKspec and continued script

HIP92177 telluric for PN_Vy2-2
  1. sky 0170 src 0171 slit 0172 BBslit 0173 confirm img
  2. 0175 zJspec
  3. 0177 HKspec

09:53 V1974Cyg
Modified script to shorten the DITs.
  1. 0197 BrG: 12x5.0s (420 ADU in 5 sec) Ks 20x2.75s (3kADU in 2.75 sec)

>>>It would be nice on the RTD to be able to set a "sticky" background image that is

automatically subtracted from subsequent images until replaced with a new background
image or the "sticky" bit is unselected and it goes back to the current paradigm

HIP101748 G210 Throughput
It is quite clear right now, take another set for the grating throughput
EL=78 and rising, IQ is again not great
  1. src 0199 slit 0200 z-only slit 0201 confirm
  2. 0227

FIRE (actually humidifier) ALARM went on. Had to clear the observing room and go to
the lobby until cleared.

11:24 HIP 87643 G210 Throughput
No time to take the spectra...
...0229

11:39 Twilight flats
No generic dither script...dither manually on telescope service

Adjusted FM4m1 -3000 steps

0230 0266 1 x 3.0 all data. Sequence offsets done by hand, probably some mistakes.
Intended was sets of 5 images doing: J H (lowflux) J H (highflux) Ks (low then high).
Will have to check headers to figure out what is what.

12:05 Steve is closing up

-- %USERSIG{JochenHeidt - 2015-05-29}%
Topic revision: r6 - 04 Aug 2015, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback